Solve
5/3 - 2/x-4 =1/2

Answer by stanbon(75887)

Solve
5/3 - 2/x-4 =1/2
---
lcm of the denominators: 6(x-4)
------------
Multiply thru by the lcm to get:
5*2*(x-4) - 2*6 = 3(x-4)
---
7(x-4) = 12
x-4 = 12/7
x = 1.7143 + 4
---
x = 5.7143
===============
Cheers,
Stan H.